Three-circle, red-on-white is one distinctive pattern painted on ceramic vessels of the Anasazi period found at an archaeological site. At one excavation, a sample of 163 potsherds indicated that 70 were of the three-circle, red-on-white pattern.

(a) Find a point estimate p̂ for the proportion of all ceramic potsherds at this site that are of the three-circle, red-on-white pattern. (Round your answer to four decimal places.)


(b) Compute a 95% confidence interval for the population proportion p of all ceramic potsherds with this distinctive pattern found at the site. (Round your answers to three decimal places.)

lower limit =

upper limit =
Answer #1

Solution :

Given that,

n = 163

x = 70

Point estimate = sample proportion = = x / n = 70 / 163 = 0.4294

1 - = 0.5706

Z/2 = 1.96

Margin of error = E = Z / 2 * (( * (1 - )) / n)

= 1.96 * (((0.4294 * 0.5706) / 163)

= 0.076

A 95% confidence interval for population proportion p is ,

- E < p < + E

0.4294 - 0.076 < p < 0.4294 + 0.076

0.353 < p < 0.505

Lower limit = 0.353

Upper limit = 0.505